OpenStudy (anonymous):

Find all solutions of the equation cosx(2sinx+1)=0 The answer is A+kpi, B+2kpi and C+2kpi where k is any integer and 0

OpenStudy (anonymous):

I've tried A=1/2pi B=3/2pi and C=11/6pi but I don't know where I went wrong !

OpenStudy (anonymous):

I don't see any problem with your solutions, they all balance the equation and fit the given parameters.

OpenStudy (anonymous):

That's what I thought but apparently one of the answers I'm giving is wrong

OpenStudy (anonymous):

7pi/6 also works

OpenStudy (anonymous):

I tried that too :/